# Hire a Linguist!: Learning Endangered Languages with In-Context Linguistic Descriptions

### v1

@@ Abstract
How can large language models (LLMs) process and translate endangered languages? Many languages lack a large corpus to train a decent LLM; therefore existing LLMs rarely perform well in unseen, endangered languages. On the contrary, we observe that 2000 endangered languages, though without a large corpus, have a grammar book or a dictionary. We propose LingoLLM, a training-free approach to enable an LLM to process unseen languages that hardly occur in its pre-training. Our key insight is to demonstrate linguistic knowledge of an unseen language in an LLM’s prompt, including a dictionary, a grammar book, and morphologically analyzed input text. We implement LingoLLM on top of two models, GPT-4 and Mixtral, and evaluate their performance on 5 tasks across endangered or low-resource languages. Our results show that LingoLLM elevates translation capability from GPT-4’s to BLEU for language directions. Our findings demonstrate the tremendous value of linguistic knowledge in the age of LLMs for endangered languages. Our data, code, and model generations can be found at https://github.com/LLiLab/llm4endangeredlang.
@@ 1 Introduction
Large language models (LLMs) are already powerful in many language understanding and generation tasks Brown et al. (2020); Ouyang et al. (2022). Their language processing capabilities rely on very large amounts of training data Kaplan et al. (2020); Hoffmann et al. (2022). For example, a recent LLM Llama-2 uses a pre-training dataset with trillion tokens (Touvron et al., 2023). While languages such as English or Spanish enjoy abundant accessible data, the majority of the world’s 7000 languages lack a rich corpus, including most endangered languages recognized by UNESCO (Moseley, 2010). Existing LLMs such as Llama Touvron et al. (2023) and GPT-4 show poor performance on languages that may not occur in pre-training Robinson et al. (2023). We believe that speakers of endangered languages deserve equitable access to NLP technologies including LLMs. How can we enable an LLM with language processing capabilities on unseen and endangered languages?
@@ 4 Experiment
Input |
suweni geren xusai dorgi de nikan i niyalma udu qoohiyan i niyalma udu |
Way ts’ax wildiihl hehl Gitwinhlguu’l ii needii hasakdiit ehl reserve. "Needii hasaga’m dim dip suwii gi’namhl laxyibi’m," dihiida. |
nihcihcee3ciiteit niiyou nuh’uuno heenees3i’okuutooni’ |
Few-Shot |
Every person in the military and every person in the common people must have courage |
He said, "I will stay here in Gitanyow, and you will go to the reserve. ’You will learn to speak English well there,’ he told me." |
I’m going to work for you tomorrow. |
| LingoLLM |
How many Chinese people and how many Koreans are there among your numerous students? |
"Although it seems that the people of Kitwancool don’t want the reserve, ’We do not wish to give away our land,’” they said. |
Someone accidentally entered this room where people sit. |
Ground Truth |
Among your many students, how many are Chinese and how many are Korean? |
And now even though the people of Kitwancool said they did not want the little reserve; "We don’t want to give away our land," they said. |
He inadvertently walked in where peope were sititng . |
@@ 4.2 Results
BLEURT |
Input Example |
Translation Example |
Dict. Only |
0.4573 |
what shoot goose moose |
Did the moose shoot the grouse? |
Dict. + Morph. Analysis |
0.5448 (+19%) |
what-CN CONTR shoot-TR-2SG grouse IRR moose |
What did you shoot over there, a grouse or a moose? |
Ground Truth |
What did you shoot? A grouse or a moose? |

### v2

# Hire a Linguist!: Learning Endangered Languages in LLMs with In-Context Linguistic Descriptions

@@ Abstract
How can large language models (LLMs) process and translate endangered languages? Many languages lack a large corpus to train a decent LLM; therefore existing LLMs rarely perform well in unseen, endangered languages. On the contrary, we observe that 2000 endangered languages, though without a large corpus, have a grammar book or a dictionary. We propose LingoLLM, a training-free approach to enable an LLM to process unseen languages that hardly occur in its pre-training. Our key insight is to demonstrate linguistic knowledge of an unseen language in an LLM’s prompt, including a dictionary, a grammar book, and morphologically analyzed input text. We implement LingoLLM on top of two models, GPT-4 and Mixtral, and evaluate their performance on 5 tasks across endangered or low-resource languages. Our results show that LingoLLM elevates translation capability from GPT-4’s to BLEU for language directions. Our findings demonstrate the tremendous value of linguistic knowledge in the age of LLMs for endangered languages. Our data, code, and model generations can be found at https://github.com/LeiLiLab/LingoLLM.
Hire a Linguist!: Learning Endangered Languages
in LLMs with In-Context Linguistic Descriptions
Kexun Zhang1 Yee Man Choi1 Zhenqiao Song1 Taiqi He1 William Yang Wang2 Lei Li1 1Carnegie Mellon University 2UC Santa Barbara {kexunz,yeemanc,zhenqias,taiqih,leili}@cs.cmu.edu william@ucsb.edu
@@ 1 Introduction
Large language models (LLMs) are already powerful in many language understanding and generation tasks Brown et al. (2020); Ouyang et al. (2022). Their language processing capabilities rely on very large amounts of training data Kaplan et al. (2020); Hoffmann et al. (2022). For example, a recent LLM Llama-2 uses a pre-training dataset with trillion tokens (Touvron et al., 2023). While languages such as English or Spanish enjoy abundant accessible data, the majority of the world’s 7000 languages lack a rich corpus, including most endangered languages recognized by UNESCO (Moseley, 2010). Existing LLMs such as Llama Touvron et al. (2023) and GPT-4 show poor performance on languages that may not occur in pre-training Robinson et al. (2023). The previous single machine translation model LegoMT Yuan et al. (2023) supports 433 languages, the largest number of languages, but it cannot extend to endangered languages. We believe that speakers of endangered languages deserve equitable access to NLP technologies including LLMs. How can we enable an LLM with language processing capabilities on unseen and endangered languages?
@@ 4 Experiment
Input | suweni geren xusai dorgi de nikan i niyalma udu qoohiyan i niyalma udu | Way ts’ax wildiihl hehl Gitwinhlguu’l ii needii hasakdiit ehl reserve. "Needii hasaga’m dim dip suwii gi’namhl laxyibi’m," dihiida. | nihcihcee3ciiteit niiyou nuh’uuno heenees3i’okuutooni’ |
Few-Shot | Every person in the military and every person in the common people must have courage | He said, "I will stay here in Gitanyow, and you will go to the reserve. ’You will learn to speak English well there,’ he told me." | I’m going to work for you tomorrow. |
LingoLLM | How many Chinese people and how many Koreans are there among your numerous students? | "Although it seems that the people of Kitwancool don’t want the reserve, ’We do not wish to give away our land,’” they said. | Someone accidentally entered this room where people sit. |
Ground Truth | Among your many students, how many are Chinese and how many are Korean? | And now even though the people of Kitwancool said they did not want the little reserve; "We don’t want to give away our land," they said. | He inadvertently walked in where peope were sititng . |
@@ 4.2 Results
BLEURT | Input Example | Translation Example | |
Dict. Only | 0.4573 | what shoot goose moose | Did the moose shoot the grouse? |
Dict. + Morph. Analysis | 0.5448 (+19%) | what-CN CONTR shoot-TR-2SG grouse IRR moose | What did you shoot over there, a grouse or a moose? |
Ground Truth | - | - | What did you shoot? A grouse or a moose? |
@@ 5 Ablation and Analysis
5.4 Comparing LingoLLM with a human baseline.
To give the readers a better idea of LingoLLM’s performance, we have asked one of the authors of this paper to mimic the behavior of LingoLLM in translating Manchu to English. They have not been exposed to our Manchu data and have no prior knowledge of Manchu. We ask them to finish the translation of 35 Manchu sentences in a limited amount of time (2 hours), which is longer than the time cost for LLMs. We report both human and model’s performance in Table 5. The superior performance of human indicates that although LLMs can already learn a lot about a new language from its grammar book and dictionary, they are still not as good as human beings on this task.
Few-Shot | LingoLLM | Human | |
BLEU | 0.82 | 9.12 | 20.32 |
5.5 LingoLLM as a benchmark for long-context understanding.
LingoLLM really requires its backbone LLM to have good long-context understanding ability as grammar books can easily exceed 50000 tokens. Not only does the LLM need a large context size, it must also be good at finding the relevant chapters and examples in the book to perform well in LingoLLM. Therefore, we can use the performance with LingoLLM as a metric for comparing LLMs. To demonstrate this possibility, we evaluate 4 backbone models on the mnc to eng task and report the results in Table 6. The performance gaps between different models are significant even for similar-sized models such as Olmo-7B and Mistral-7B.
Olmo-7B | Mistral-7B | Mixtral-8x7B | GPT-4 |
0.54 | 3.71 | 4.4 | 10.8 |
@@ 7 Conclusion
Acknowledgements
L.L. is partly supported by a gift from Apple Inc.
We are grateful to Lori Levin for their suggestions and helpful discussions. Thanks also to Danqing Wang and Yuanjing Wei for their proof-reading and comments. We appreciate the reviewers of this paper for their engagement in the review process. Special thanks to the native Manchu speakers who provided valuable insights on solving math problems in Manchu and annotated the math reasoning benchmark for us.
@@ Appendix C Language and Their ISO 639-3 Code
Appendix D More translation examples
Ground Truth | LingoLLM | BLEURT | Rank |
Now where are you going? | Where are you going now? | 0.86 | 1% |
Among your many classmates, how many are Chinese and how many are Korean? | How many Chinese people and how many Koreans are there among your numerous students? | 0.74 | 5% |
A letter from home is worth ten thousand liang of gold. | The letter of the house is worth ten thousand ounces of gold. | 0.69 | 11% |
I’m going to Beijing (the capital city) | I am going toward the city’s capital. | 0.65 | 20% |
What books does he explain? | What book are we explaining? | 0.62 | 30% |
Don’t worry for us, it’s no big deal. | You need not be very distressed on our account, it doesn’t matter. | 0.59 | 40% |
I live in Liaodong city. | I have resided in the inner part of the walled city of Liaodong. | 0.53 | 50% |
Master, light a lamp and bring it | Having lit the lamp, bring it (along), master. | 0.50 | 60% |
I’m a Korean person, I don’t walk with familiarity in places of China | I, even if a person of Korea, was not sad about walking the land of the Chinese. | 0.47 | 70% |
Why do you learn Chinese language? | You, teach that Chinese person’s book, how? | 0.44 | 80% |
This is very well, we should go together! | If so, we will likely do it together, the partridge. | 0.40 | 90% |
Appendix E Math reasoning examples
Question in Manchu:
Balder hvqin deri muke tunggiyeme gaimbi. erde sunja hunio, yamji ninggun hunio tunggiyembi. emu hunio de sunja moro biqi, tere inenggi moro udu muke bahambi?
Question in English (Ground Truth):
Baldur gets water from a well. He gets 5 pails of water every morning and 6 pails of water every afternoon. If each pail contains 5 liters of water, how many liters of water does he get every day?
Question in English (LingoLLM):
Balder, early in the morning, picks up water from the well. He takes five buckets in the evening, and six buckets in the morning. If one bucket equals five bowls, how many bowls of water does he get in a day?
Verbatim Output of LingoLLM:
Question in Manchu: gvsin ninggun (36) penguin ni nimanggi tala de xun fiyakamahabi. baji ome ilan i emken (1/3) mederi qi dosifi ebxeme genefi, jai ilan i emken (1/3) dunggu de jaka jeme genefi. geli ya labdu penguin kemuni xun fiyakame ilihabi?
Question in English (verbatim): There are 36 penguins sunbathing in the snow. One-third of them jump in and swim in the ocean. Another one-third go inside the cave to eat their dinner. How many penguins are still left sunbathing?
Question in English (LingoLLM): Thirty-six penguins on the snowy plain sled every day. A little more than one-third of them hurriedly entered the sea, and another one-third went into the cave to eat something. How many penguins are still sledding every day?